OpenStudy (anonymous):

I'm trying to find out how many grams of water can be produced in a reaction of 11.7 moles of ethane with oxygen gas. The balanced equation is 2C2H6 + 7O2 yields 4CO2 + 6H2O. I need help determining the setup of the conversion. I think it would be 11.7 mol C2H6 x 54.039 mol H2O/1 mol C2H6 x something over 3 mol H2O...help? :)

OpenStudy (anonymous):

@Australopithecus

OpenStudy (anonymous):

@waxynaveed

OpenStudy (anonymous):

from balanced equation, 2 moles of ethane can produce= 6 moles of water 11.7 moles . .. .. .. .. . . . .. . = 6/2 X 11.7 ...................................................... = 35.1 moles of water mass of water = moles X m.mass of H2O mass = 35.1 X 18 = 631.8 g of water will produce when 11.7 moles of ethane will react with oxygen

OpenStudy (australopithecus):

So this is simple, 2C2H6 + 7O2 -> 4CO2 + 6H2O First off your limiting reagent is the ethanol 11.7 mole you assume oxygen in excess in this question so for every 2 moles of ethanol reacted 6 moles of water are produced that is what the equation is telling me, so divide by 2, then multiply by 6 to find the moles of water produced, then simply use the equation, moles = grams/molecular mass in this case you want the molecular mass of water

OpenStudy (anonymous):

Where did you get the 18 from???

OpenStudy (australopithecus):

18 is the molecular mass of water, he should have put units grams/mole

OpenStudy (australopithecus):

15.9994g/mol for the molecular mass of oxygen 1.0079g/mol for the molecular mass of hydrogen he just rounded them, so, H2O consists of two hydrogens and one oxygen so, 1.0g/mol + 1.0g/mol + 16g/mol = 18g/mol molecular mass

OpenStudy (anonymous):

H2O= 2+16= 18 molar mass of water

OpenStudy (australopithecus):

roughly the molecular mass of water